40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 129     Entry time: Mon May 24 19:40:12 2010
Author: tarac 
Type: Laser 
Category: Laser 
Subject: Ref Cav prealignment 

 

Dear elog

  I don't know why I can't lock the PMC. After changing the 21.5 MHz card because of the loosen SMA connector, the 21.5 MHz EOM is working.

The error signal looks good. I adjusted the gain, flipped the phase by 180 degree, and still cannot lock the cavity. 

(The medm was frozen this morning, Peter helped reset it back to work this after noon.) It might be insufficient amount of power coupling into the cavity.

The minimum reflected beam I could get is only ~1/2 of the total DC power, I'll try to align the beam and move the lens a little bit more to see if I can optimize it better.

So I skip this PMC part for now, and pre align the path to mode cleaner. The mode matching is ok. Two lens, f1=114.5 and f2=286.3 mm are good. A CCD behind the cavity is set in place.

 

ELOG V3.1.3-